# Hype-Adjusted Probability Measure for NLP Stock Return Forecasting

### v2

@@ Abstract
This manuscript introduces the Hype-Adjusted Probability Measure developed in the context of a new Natural Language Processing (NLP) approach for stock return and volatility forecasting. A novel sentiment score equation is presented to capture component and memory effects and assign dynamic parameters, enhancing the impact of intraday news data on forecasting next-period volatility for selected U.S. semiconductor tickers. This approach integrates machine learning techniques to analyze and improve the predictive value of news. Building on the research of Geman et al [deveikyte2022], this work improves forecast accuracy by addressing news bias, memory, and weight, and incorporating shifts in sentiment direction. Finally, we propose the Hype-Adjusted Probability Measure, proving its existence and uniqueness, and discuss its theoretical applications in finance for NLP-based stock return forecasting, outlining future research pathways inspired by its concepts.
@@ 1 Introduction
Attention generates capital; capital shifts draw attention, and hype serves as the catalyst.
The inherent randomness of financial markets implies that precise predictions of future volatility and returns for any asset or its derivatives are almost surely impossible. Many pioneers have built sophisticated analytical and numerical instruments to evaluate future prices under strict assumptions, such as the well-known Black-Scholes model for an ideal risk-neutral world [BlackScholes1973]. Extensive research has shown that calibration and theoretical assumptions are essential to enhance model performance.
Sentiment has long been recognized as a key driver in financial markets, influencing both trading behavior and asset prices. Previous studies have examined the theoretical and empirical role of sentiment in market dynamics, highlighting its impact on investors’ decision-making and trading patterns [hirshleifer2001role, zhong2018sentiment]. Furthermore, research shows a significant relationship between investor sentiment and stock returns, providing a foundation for sentiment-based trading strategies [baker2006investor].
This manuscript contains key observations from “A Sentiment Analysis Approach to the Prediction of Market Volatility", where Latent Dirichlet Allocation (LDA) is applied to forecast market price and volatility [deveikyte2022]. The authors suggested an advanced method to compute the sentiment score and introduced how to forecast the magnitudes of the next day’s market return and volatility, in addition to predicting future stock return and volatility change.
We choose semiconductors as the primary sector for analyzing news and market data. We first review previous works, and propose the market intuition of the “3 weights" and numerical improvements based on the new sentiment score equation and NLP approach through the improved forecast results using ML.
We first review the past literature on NLP and Change of Measure in section 2. A detailed overview of data and methods is presented in section 3 which includes the developments from previous sentiment equation to more sophisticated ones along with enhanced NLP forecasting methods. Building upon these foundations, we introduce a novel probability measure, the Hype-Adjusted Probability Measure, $\mathbb{P}^{a}$, in section 4, designed to capture the effects of market "hype." Lastly, the results and future work are discussed in section 5.
@@ 2 Literature Review
2.1 Literature Review for NLP
Sentiment analysis, a subfield of Natural Language Processing (NLP), focuses on quantifying the emotional tone and intent conveyed in textual data. NLP itself is a rapidly growing area of machine learning (ML) that enables computers to process and understand human language through algorithms and statistical models [jurafsky2009speech]. Within NLP, sentiment analysis applies techniques to assess positive, negative, or neutral sentiment in data, often used for applications like market forecasting and consumer behavior analysis [liu2012sentiment].
In this context, sentiment is defined as the emotional polarity associated with a piece of text, often derived from news articles, tweets, or other media. It is typically measured using computational methods such as dictionary-based approaches or machine learning models. For example, Vader, a lexicon-based sentiment analysis tool, evaluates sentiment scores ranging from extremely negative to extremely positive [hutto2014vader].
In their previous research, Hélyette Geman and her colleagues identified a significant negative correlation between daily negative sentiment and observed market volatility and developed a sentiment-based model for predicting directional volatility and market returns [deveikyte2022]. Building on their findings, we aimed to enhance the predictive relationship between news sentiment and stock performance by refining the methods used to calculate key financial metrics.
The return on day $t$ is defined as the logarithmic change in the closing price from day $t-1$, expressed as:
@@ 2.1 Literature Review for NLP
Additionally, we adopted their method for defining the volatility, $\sigma$, which is calculated as:
@@
Notice, for the programming and machine learning components of the manuscript, we selected a rolling window of 5 days to account for the amount of a regular trading week (without holiday breaks).
@@
The base sentiment function introduced in [deveikyte2022] is defined as the fraction of the difference between the number of positive and negative sentiment tokens, scaled by the sum of positive, neutral, and adjusted negative cases:
@@
One significant limitation of the prior sentiment equation is that it only counts the number of positive, negative, and neutral news. A news of sentiment score $-3$ should carry a bigger weight than a $-0.1$ score.
As a part of his master’s thesis, Zheng Cao, lead author of this paper, proposed a modified approach to sentiment scoring in April 2024 at Johns Hopkins University [CaoGemanThesis]. First, an interval of $(-0.05,+0.05)$ was chosen to broaden the range of sentiment scores classified as neutral, setting all scores within this domain to 00. Additionally, sentiment scores of extreme values, such as +4, are given higher weights than minimal positive values (e.g., +0.1), addressing the limitation in the original model where they had equal weights.
$Sent_{d}=\frac{1}{n}\sum_{i=1}^{n}SentScore_{i,d}$ (4)
Cao further incorporated a time-memory weight effect in sentiment calculations, represented as follows:
$Sent_{d}=$ $\omega_{\text{time weight},d}\frac{1}{n}\sum_{i=1}^{n}SentScore_{i,d}$ (5)
$+\omega_{\text{time weight},d-1}Sent_{d-1}$ (6)
These equations serve as the foundation for the developments of more sophisticated sentiment equations and the proposal of the hype-adjusted probability measure in the later sections.
2.2 Literature Review for Change of Measure
We draw inspiration for the change of measure from seminal works such as Harrison and Kreps’ 1979 introduction of the risk-adjusted measure for option pricing, which laid the foundation for modern financial mathematics by connecting the physical measure to pricing probabilities [HarrisonKrep1979]. Their work established the most fundamental cornerstone in this field, focusing primarily on stock returns rather than option pricing.
Building on this, Geman’s groundbreaking work in her 1989 dissertation, later published in “Changes of Numéraire, Changes of Probability Measure and Option Pricing" [Geman1995], extended these ideas to option pricing with stochastic interest rates by showing that changes in probability measures are intrinsically linked to changes in the Nnméraire.
More recently, Glasserman, Mamaysky, and Qin (2023) introduced the concept of entropy-based news novelty as a predictor of negative stock market returns and macroeconomic outcomes [glasserman2023new]. Their work leverages changes in the distribution of news text, quantified through entropy measures derived from recurrent neural networks. This innovative approach highlights the role of shifting language patterns in financial forecasting, offering a complementary perspective to traditional probability measure transformations.
In our work, while we acknowledge these advancements, we focus specifically on the representation of stock returns under measure changes. Although Geman et al.’s remarkable insights on measure changes for option pricing provide a rich theoretical basis, we intentionally leave the detailed discussion of these concepts for future research. The present paper aims to introduce the change of measure framework in the context of stock returns and sentiment-driven probability measures, laying the groundwork for further exploration in option pricing and beyond.
@@ 3 Data and Methods
In this study, ‘news’ refers to textual data from verified financial sources reporting on market trends, corporate developments, and stock performance.
Primary data sources for this study were obtained from LSEG and the Eikon API [LSEG], which provided access to over half a million news articles related to 30 semiconductor stocks. These data sources offered comprehensive coverage of market events, corporate developments, and sentiment indicators.
@@
    Assigning Weights to News Sources: Each news source was assigned a weight to achieve a more objective sentiment score adjustment. The adjusted sentiment score was calculated as:
@@
    Adjusting for Over- or Under-Reported Events: To address the uneven representation of events in the dataset, we applied two key measures:
    • Removed duplicate or near-identical articles to prevent overrepresentation of certain events.
    • Amplified the impact of under-reported events by assigning higher weights or using clustering techniques to identify and include diverse event representations.
3.1 Data and Naive Model Results
Appendix A presents a sample table of 30 tickers selected from the iShares Semiconductor ETF (SOXX). Note This ticker weight table is adjusted by removing 2 values: CME E-MINI S&P500-TECHNOLOGY SECTOR INDEX FUTURE SEP 2024,$0.15247\%$ and CME INDEX and OPTIONS MARKET E-MINI RUSSELL 2000 INDEX FUTURE SEP 2024, $0.15247\%$. The reason is that these two entries do not represent any actual company; they are financial derivatives or futures contracts tied to broader indices rather than individual equities, and therefore their inclusion would distort the sector-specific analysis focused on individual component stocks.
@@ 3.1 Data and Naive Model Results
Each data set generates a confusion matrix report through a simple Linear Discriminant Analysis algorithm [balakrishnama1998lda]. Linear Discriminant Analysis (LDA) is a commonly used dimensionality reduction technique that is effective for classification. It projects data onto a lower-dimensional space while maximizing separation between classes by modeling inter-class differences using the mean and variance within each class. Unlike Principal Component Analysis (PCA), which maximizes variance, LDA prioritizes class separability, making it especially useful in supervised learning with labeled data.
@@
The latter sections present a more sophisticated equation of processing the sentiment score, with supporting results provided.
@@ 3.2.1 News Bias
  • Under-represented tickers may not sufficiently contribute to the overall sentiment, resulting in overlooked trends.
@@
We examine news bias within the iShares Semiconductor ETF (SOXX) of the U.S. stock market by analyzing how individual asset components contribute to the overall index. Several major companies, such as Nvidia, dominate the sector, with Nvidia representing over $8\%$ of the market. However, over the past 15 months (up to mid-July 2024), Nvidia has received disproportionately high news coverage of $24.52\%$ compared to other companies in the sector. Considering 30 companies from the SOXX, we assign weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%. Please refer to Appendix A for more details.
@@
We wish to provide some sophisticated method to change the relationship from the first plot Figure 2 to the second in Figure 4.
@@
Further data and analysis are provided in the Intuition of the New Measure subsection.
To address news bias, we consider two key factors:
  1. 1.
    Event Exaggeration: Some media outlets tend to sensationalize events, which can disproportionately impact investor decisions and market reactions.
  2. 2.
    Selective Focus on High-Profile Companies: Media often emphasizes a few prominent companies, amplifying their perceived influence and neglecting others.
We propose two approaches to address these biases:
  1. 1.
    Assigning Weights to News Sources: Each news source is assigned a weight to adjust sentiment scores more objectively. For example:
    $\text{sentiment}_{adj}=\alpha\cdot\text{sentiment}+\beta$ (10)
    where $\alpha$ controls the sentiment weight from the source, and $\beta$ corrects for inherent bias.
  2. 2.
    Adjusting for over- or under-reporting: To correct for uneven media coverage, we adjust the volume of reported news by removing duplicates or overly similar articles.
@@ 3.2.2 Memory and Weight
Quantifying the bias of news sources and optimizing compound sentiment scores are key to improving sentiment-based forecasting models. We propose a weighted approach that integrates current sentiment with past sentiment, adjusting for biases:
$Sent_{d}=$ $\ \omega_{\text{today weight}}\sum_{i=1}^{n}\omega_{\text{news bias}}\omega_{\text{component weight}}\,SentScore_{i,d}$ (11)
$+\omega_{\text{past weight}}Sent_{d-1}$ (12)
@@
  • $d$: Represents today or the current period of time. $d-1$ refers to yesterday or the previous period of time.
@@
  • Present sentiment: A weighted sum of sentiment scores adjusted for news bias ($\omega_{\text{news bias}}$) and component weights ($\omega_{\text{component weight}}$).
@@
Future work can be to continue improving the quantifying method for the news bias.
3.2.3 Recursive Method
We propose new parameters for calculating the daily compound sentiment score. Rather than using daily averages as a benchmark for $Sent_{d}$ (sentiment score on the date or time $d$), we assign weights to each component (ticker) based on its sector weight:
$+\omega_{\text{past weight}}Sent_{d-1}$ (14)
Using the semiconductor sector as an example, we analyze how individual asset components contribute to the overall index. Considering 30 companies from the SOXX, we assign weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%.
$\sum_{n=0}^{29}\omega_{n}=1$ (15)
We introduce a novel approach to account for time-memory effects in market reactions. Historical news sentiment impacts future market performance, with older news exerting diminishing influence. For instance, the collective sentiment on a Monday may be shaped not only by weekend news but also by earlier events. We propose weighting algorithms to capture these decay and lagging effects, ensuring older events carry less influence while accounting for delayed impacts across different assets.
@@ 2. News Source Bias Weights:
We recognize that news reports are rarely perfectly objective. Different news sources have biases that can affect sentiment. For example, CNN might exhibit a liberal bias and present a relatively positive outlook on green energy, whereas FOX may be more critical of gun control policies. Similarly, the BBC may adopt a more critical stance on the Chinese market compared to state-owned Chinese media, with biases shaped by the source’s audience and language.
@@ 3. Weights of Past News Data (Memory):
We have already demonstrated that adjusting sentiment using indicator functions, coupled with machine learning techniques to optimize parameters, can enhance the accuracy of market forecasts.
@@
In summary, we focus on evaluating and optimizing the “3 weights" to enhance the sentiment score equation and NLP approach. These enhancements ensure a balanced and accurate representation of market dynamics by correcting for media attention imbalances, accounting for source biases, and incorporating memory effects. This foundational framework not only refines sentiment modeling but also provides the necessary groundwork for the development of an improved sentiment score equation. In the next section, we extend these concepts to introduce a more robust equation that integrates component weighting and historical memory effects for greater predictive accuracy.
3.3 Improved Sentiment Score Equation
This section presents an enhanced sentiment score equation that incorporates the weighting of individual components within a portfolio and accounts for historical memory effects from past news data.
Notations
  1. $1$
  2. $2$
    $SentScore_{i,d}$: The average sentiment score for news related to ticker $i$ on date $d$, based on a total of $k$ news articles.
  3. $3$
$Sent_{d}=$ $\ \sum_{i=1}^{n}\omega_{i}\,SentScore_{i,d}$ (16)
where $n$ represents the total number of tickers in the portfolio.
$SentScore_{i,d}=\frac{1}{k}\ \sum_{j=1}^{k}S_{i,d,j}$ (17)
where $SentScore_{i,d}$ denotes the average sentiment score for ticker $i$ on date $d$, calculated from a total of $k$ news articles.

### v4

# A Hype-Adjusted Probability Measure for NLP Stock Return Forecasting

@@ Abstract
This article introduces a Hype-Adjusted Probability Measure in the context of a new Natural Language Processing (NLP) approach for stock return and volatility forecasting. A novel sentiment score equation is proposed to represent the impact of intraday news on forecasting next-period stock return and volatility for selected U.S. semiconductor tickers, a very vibrant industry sector. This work improves the forecast accuracy by addressing news bias, memory, and weight, and incorporating shifts in sentiment direction. More importantly, it extends the use of the remarkable tool of change of Probability Measure developed in the finance of Asset Pricing to NLP forecasting by constructing a Hype-Adjusted Probability Measure, obtained from a redistribution of the weights in the probability space, meant to correct for excessive or insufficient news.
@@ 1 Introduction
Sentiment has long been recognized as a key driver in financial markets, influencing both trading behavior and asset prices. Previous studies by Hirshleifer (2001) and Zhong (2018) have examined the theoretical and empirical role of sentiment in market dynamics, highlighting its impact on investors’ decision-making and trading patterns[hirshleifer2001role, zhong2018sentiment]. Furthermore, research [baker2006investor] by Baker and Wurgler (2006) followed by others has shown a significant relationship between investor sentiment and stock returns, providing a foundation for sentiment-based trading strategies.
The NLP part of our manuscript draws inspiration from the work of Deveikyte et al (2022) [deveikyte2022], who applied Latent Dirichlet Allocation to forecast market prices and volatility. The authors proposed a thorough approach to compute the sentiment score and forecast the directions of the next day’s market return and volatility in the case of FTSE100 stocks. In this research, we choose semiconductors, a very vibrant industry of this decade, as the sector for analyzing news and market data.
After a review of the literature on NLP and Changes of Measure in section 2, a detailed overview of data and methodology is presented in section 3, which includes the developments from previous sentiment equation to more sophisticated ones along with enhanced NLP forecasting methods. Building upon these foundations, we introduce in section 4, a novel probability measure, that we name “ Hype-Adjusted Probability Measure", designed to capture the occurrence of market “hype." The results and future work are discussed in section 5, the forecasting accuracy is significantly increased by $8\%$.
To our best knowledge, no paper has yet introduced a change of probability measure in the context of NLP forecasting of asset returns and volatility. The instrument is beautiful, the application here original.
@@ 2 Literature Review
Sentiment analysis, a subfield of Natural Language Processing (NLP), focuses on quantifying the emotional tone and intent conveyed in textual data. NLP itself is a rapidly growing area of machine learning (ML) that enables computers to process and understand human language through algorithms and statistical models such as the founding work by Jurafsky and Martin (2000) [jurafsky2009speech]. Within NLP, sentiment analysis applies techniques to assess positive, negative, or neutral sentiment in data, often used for applications like market forecasting and consumer behavior analysis [liu2012sentiment].
In this context, sentiment is defined as the emotional polarity associated with a piece of text, often derived from news articles, tweets, or other media. It is typically measured using computational methods such as dictionary-based approaches or ML models. For example, Vader, a lexicon-based sentiment analysis tool, evaluates sentiment scores ranging from extremely negative to extremely positive [hutto2014vader].
In their 2022 paper, Deveikyte et al identified a significant negative correlation between daily negative sentiment and observed market volatility and develop a sentiment-based model for predicting directional volatility and market returns [deveikyte2022]. Building on their findings, we aimed to enhance the predictive relationship between news sentiment and stock performance by refining the methods used to calculate key financial metrics.
The return on day $t$ are classically defined as the logarithmic change in the closing price from day $t-1$, expressed as:
@@
The annualized volatility, $\sigma$, is calculated by the formula:
@@
Notice, for the programming and ML components of the manuscript, we selected a rolling window of 5 days to account for the amount of a regular trading week (without holiday breaks).
@@
The sentiment score has been defined by many authors, in particular the paper [gabrovsek2016twitter] as the fraction of the difference between the number $N$ of positive and negative sentiment tokens, scaled by the sum of positive, neutral, and adjusted negative cases, of a given time $d$:
@@
Glasserman and Mamaysky (2017) highlight the predictive power of unusual news patterns in forecasting market stress [glasserman2017unusualnews]. Shapiro et al (2022) developed a novel sentiment-scoring model tailored to economic news articles, showing that daily news sentiment predicts consumer sentiment movements and macroeconomic responses to sentiment shocks, such as increased consumption, output, and interest rates [shapiro2020measuring]. Cohen et al. (2023) demonstrate the robustness of multimodal classifiers under perturbation attacks and present the potential for integrating resilient, multimodal approaches into financial sentiment analysis for improved market forecasting [cohen2023masking].
The change of probability measures has been revealed as a remarkable tool in the finance of Asset Pricing. It started with Paul Samuelson’s “Proof that properly anticipated prices fluctuate randomly" (1965) that he identified the importance of change of probability measure in Finance and prices of futures contracts are martingales [samuelson1965]. Harrison and Kreps (1979) proposed under No Arbitrage and constant interest rates a new risk-neutral (equivalent martingale) probability measure that incorporates the risk premium attached to equities and leads to martingales for discounted stock prices[HarrisonKrep1979]. The “forward measure" introduced in Geman (1989) [Geman1989] was constructed to address the challenges created by stochastic interest rates in the valuation of risky cash flows.
To recall, given a random variable $X$ and a filtration $\mathcal{F}$, the conditional expectation $\mathbb{E}[X\mid\mathcal{F}]$ is defined as:
[…]
where $\mathcal{F}$ represents the information up to the current time, and $\Omega$ is the set of states of nature. This concept underpins the mathematical framework of the Radon-Nikodym derivative, used to construct new probability measures like the risk-neutral or hype-adjusted measures in financial applications.
Note that in the context of product management, Wind and Mahajan (1987) emphasized the value of “marketing hype" when launching new products in order to create a supportive market environment.
@@ 3 Data and Methods
As in the foundational paper of Bengio et al (2003) [bengio2003neural], we define in this paper ‘news’ as textual data, extracted from verified financial sources reporting on market trends, corporate developments, and stock performance. Primary data sources for this study were obtained from LSEG and the Eikon API [LSEG], which provide access to over half a million news articles related to 30 semiconductor stocks. These data sources offer a comprehensive coverage of market events, corporate news, and sentiment indicators.
@@
    Assigning Weights to News Sources: Each news source is assigned a weight to achieve a more objective sentiment score adjustment and the adjusted score is calculated as:
@@
    Adjusting for Over- or Under-Reported Events: To address the uneven representation of events in the dataset, we use two key methods:
    • Removal of duplicate or near-identical articles to prevent overrepresentation of certain events.
    • Increase of the impact of under-reported events by assigning higher weights explained below.
3.1 Data and LDA Model Results
Appendix A presents a sample table of 30 tickers selected from the iShares Semiconductor ETF (SOXX). Note this ticker weight table is adjusted by removing 2 values: CME E-MINI S&P500-technology sector index future for September 2024, $0.15247\%$ and CME Index and Options Market E-MINI Russell 2000 index future for September 2024, $0.15247\%$. The reason is that these two entries do not represent any actual company, but a mix of companies that are already in our list.
@@ 3.1 Data and LDA Model Results
Each data set generates a accuracy precision report through a simple Linear Discriminant Analysis algorithm [balakrishnama1998lda]. Linear Discriminant Analysis (LDA) is a commonly used dimensionality reduction technique that is effective for classification. It projects data onto a lower-dimensional space while maximizing separation between classes by modeling inter-class differences using the mean and variance within each class. Unlike Principal Component Analysis (PCA), which maximizes variance, LDA prioritizes class separability, making it especially useful in supervised learning with labeled data.
Tables 1, 2, and 3 below present the forecast results based on the simple LDA model, with volatility direction as the predicted target and sentiment scores as the only input training parameters.
@@
The latter sections present a more sophisticated method of processing the sentiment score, with supporting results provided.
@@ 3.2 Bias, Memory, and Weights
One significant limitation of the prior sentiment equation 3 is that it only counts the number of positive, negative, and neutral news. A news of sentiment score $-3$ should carry a bigger weight than a $-0.1$ score.
We propose a modified approach to sentiment scoring. First, an interval of $(-0.05,+0.05)$ is chosen to broaden the range of sentiment scores classified as neutral, setting all scores within this domain to 00. Additionally, sentiment scores of extreme values, such as +4, are given higher weights than minimal positive values (e.g., +0.1), addressing the limitation in the original model where they had equal weights.
$Sent_{d}=\frac{1}{n}\sum_{i=1}^{n}SentScore_{i,d}$ (6)
Notations
  1. $1$
    $n$: The total number of tickers in the portfolio.
  2. $2$
  3. $3$
  4. $4$
    $SentScore_{i,d}$: The average sentiment score for ticker $i$ on date $d$, calculated from a total of $k$ news articles’ sentiments $S_{i,d,j}$.
$SentScore_{i,d}=\frac{1}{k}\ \sum_{j=1}^{k}S_{i,d,j}$ (7)
@@ 3.2.1 News Bias
  • Under-represented tickers may not properly contribute to the overall sentiment.
@@
We examine news bias within the iShares Semiconductor ETF (SOXX) of the U.S. stock market by analyzing how individual asset components contribute to the overall index. Several major companies, such as Nvidia, dominate the sector, with Nvidia representing over $8\%$ of the market. However, over the previous 15 months (up to mid-July 2024), Nvidia had received disproportionately high news coverage of $24.52\%$ compared to other companies in the sector. Considering 30 companies from the SOXX, we assigned weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%. Please refer to Appendix A for more details.
@@
We wish to provide a sophisticated method to change the relationship from the first plot in Figure 2 to the second in Figure 4.
@@ 3.2.2 Memory and Weight
Quantifying the bias of news sources and optimizing compound sentiment scores are key to improving sentiment-based forecasting models. We propose new parameters for calculating the daily compound sentiment score. Rather than using daily averages in equation 7 as a benchmark for $Sent_{d}$ (sentiment score on the date or time $d$), we assign weights to each component (ticker) based on its sector weight:
$Sent_{d}=$ $\ \sum_{i=1}^{n}\omega_{i}\,SentScore_{i,d}$ (10)
@@
  • $d$: represents the current day.
@@
We introduce a novel approach to account for time-memory effects in market reactions. Historical news sentiment impacts future market performance, with older news exerting diminishing influence. For instance, the collective sentiment on a Monday may be shaped not only by weekend news but also by earlier events. We propose weighting algorithms to capture these decay and lagging effects, ensuring older events carry less influence but some persistence.
$+\omega_{\text{past weight}}Sent_{d-1}$
@@
  • Present sentiment: A weighted sum of sentiment scores adjusted for news bias and component weights ($\omega_{\text{component weight}}$).
@@
Following equation 11 and using the semiconductor sector as an example, we analyze how individual asset components contribute to the overall index. Considering 30 companies from the SOXX, we assign weights to each ticker based on their market share. For example, Nvidia (NVDA), valued at \$1,120,866.62 million and accounting for 8.64% of the sector, has a component weight of 8.64%.
$\sum_{i=0}^{n=29}\omega_{i}=1$ (12)
@@ 2. News Source Bias Weights:
We recognize that news reports are rarely perfectly objective. Different news sources have biases that can affect sentiment. For example, CNN may exhibit a liberal bias and present a relatively positive outlook on green energy, whereas FOX may be more critical of gun control policies. Similarly, the BBC may adopt a more critical stance on the Chinese market compared to state-owned Chinese media, with biases shaped by the source’s audience and language.
@@ 3. Weights of Past News Data (Memory):
We have already demonstrated that adjusting sentiment using indicator functions, coupled with ML techniques to optimize parameters, can enhance the accuracy of market forecasts.
@@
In summary, we focus on evaluating and optimizing the 3 criteria to enhance the sentiment score equation and NLP approach. These enhancements ensure a balanced and accurate representation of market dynamics by correcting for media attention imbalances, accounting for source biases, and incorporating memory effects. This foundational framework not only refines sentiment modeling but also provides the necessary groundwork for the development of an adjusted sentiment score equation. In the next section, we extend these concepts to introduce a more robust equation that integrates component weighting and historical memory effects for greater predictive accuracy.
3.3 Adjusted Sentiment Score Equation
This section presents an enhanced sentiment score calculated based on equation 11 that incorporates the weighting of individual components within a portfolio and accounts for historical memory effects from past news data.